# One Patch to Caption Them All: A Unified Zero-Shot Captioning Framework

### v2

@@ 5 Conclusions
We introduced Patch-ioner, a novel zero-shot captioning framework that shifts from an image-centric to a patch-centric approach, enabling caption generation for individual patches and arbitrary aggregations without any region or even image supervision. We rely on the strong spatial awareness of DINOv2, whose local image patches have been effectively bridged with the text modality. Thanks to the disentangled training of the decoder network, this flexible and scalable method sets the new state of the art on various regional captioning tasks, including dense and region-based captioning, as well as our newly proposed trace captioning.
Despite its simplicity, results show that our patch-centric approach can effectively bridge the gap between local and global understanding in image captioning, providing a unified framework for multi-granularity captioning tasks in a zero-shot setting. Moreover, our models require a single backbone forward pass to caption multiple regions, facilitating practical viability in interactive applications.
@@
Despite strong zero-shot performance, our model still lags behind fully supervised, task-specific approaches. The contextual scope of each patch is fixed by the backbone and is not adjusted to meet the user intents. On top of that, the modality jump introduces noise that can cause hallucinations. Future work could incorporate weak supervision, e.g., image-level captioning loss, to improve patch-level semantics in contrastively-learned representations, or refine the patch-to-text projection to further reduce the modality gap in zero-shot settings.
Acknowledgments
This work was partially supported by the following projects: SUN – Social and hUman ceNtered XR (101092612), NextGenerationEU (FAIR PE00000013, ITSERR B53C22001770006 and MUCES B53D23026090001).
@@ 6 Implementation Details
For the training of the textual decoder of the memory-based configuration $\phi$, we adopt a prefix GPT2-style decoder-only Transformer with 4 attention heads and 4 layers, following the architecture used by Li et al. (2023b). We train the model on captions from the COCO training set, which also serves as the memory bank $M$ for the projection mechanism, comprising approximately 500k texts. We set the hyperparameters of the projection mechanism as in DeCap ($\tau=0.01$), and use the AdamW optimizer with a weight decay of 0.01. Training proceeds for 10 epochs with a learning rate of $10^{-5}$ and a batch size of 64. A comprehensive overview of the framework operating in the DeCap setting (with the projection mechanism as modality gap mitigation strategy) is provided in Figure 3.
For the external knowledge-based captioning models we performed a training of 15 epochs with a batch size of 80 captions on the same GPT2-style textual decoder using a learning rate of $2\times 10^{-5}$ and a gaussian noise variance of $16\times 10^{-3}$ to replicate the experimental settings of Fei et al. (2023) and Zeng et al. (2024).
@@ 7 Backbone Details
  • CLIP (Radford et al., 2021): A foundational model that learns a shared embedding space for images and text through contrastive learning. While being the most used model for global image-text alignment, its patch tokens are known to lack strong spatial and fine-grained semantic information. The input resolution of its training is 224 pixel.
@@
  • DenseCLIP (Rao et al., 2022): A fine-tuned version of CLIP that incorporates a pixel-text matching loss to enhance the model’s ability to understand local regions. The official implementation input resolution is 640 pixel for the ViT-B/16 version.
@@
  • INViTE (Chen et al., 2024): This method modifies CLIP’s vision transformer to bring patch tokens in the text space by disabling the self-attention mechanism. It employs the same visual encoder of CLIP, trained at 224 pixel input resolution.
@@
  • ProxyCLIP (Lan et al., 2024): A model that leverages the local understanding of a DINO backbone to improve CLIP’s patch-level representations. It achieves this by replacing the attention maps in CLIP’s final layer with DINO’s attention maps, effectively transferring DINO’s fine-grained spatial awareness to the CLIP embedding space. The DINO ViT-B/8 version was tested with images at 296 pixel resolution, while the DINOv2 ViT-B/14 at 518 pixel.
@@
  • DINO.txt (Jose et al., 2025): This model builds upon a frozen DINOv2 backbone, adding learnable transformer blocks on top. It is then trained with a contrastive objective against a text encoder to align both global and patch-level representations with language. The DINOv2 backbone was trained at the resolution of 518 pixel.
@@
  • Talk2DINO (Barsellotti et al., 2024): This model creates a bridge between the CLIP and DINOv2 embedding spaces. It trains a projection to map CLIP text embeddings into the DINOv2 patch space, using DINOv2’s highly meaningful attention maps to identify and align with the most relevant patches during training. The DINOv2 backbone was trained at the resolution of 518 pixel.
@@ 8 Additional Results: Aggregation Strategies, Input Resolution, Text Collection
Figure 3: Memory-based Patch-level captioning. Given an input image, we first extract dense patch-level representations using a vision transformer backbone. For a selected patch, we apply the projection-based mechanism introduced by Li et al. (2023b) to mitigate the modality gap and align its representation with the text embedding space. Finally, the transformed embedding is fed into a text decoder trained on a text-only corpus, generating a zero-shot caption for the patch.
@@
For patch-based captioning, we followed Talk2DINO Barsellotti et al. (2024) and used an input image resolution of 518x518, obtaining 37 14x14 patches per side when using the Talk2DINO backbone. The original DeCap Li et al. (2023b), ViECap Fei et al. (2023), MeaCap Zeng et al. (2024) implementations uses the CLIP B/32 backbone with 224x224 input images with 7 patches per side. We also tested with the CLIP B/16 backbone, resulting in 14 patches per side at 224x224 resolution, and with 592x592 input image size, to obtain the same number of patches as in our framework (37 per side).
@@
In Table 6, we report the results of standard zero-shot image captioning. In addition to the already described weighting schemes, we test one additional configuration for our framework that is central patch, where the decoding is applied to the central patch of the image. We can observe that the most effective strategy for the image captioning task is attention. This is coherent with results from Barsellotti et al. (2024), where they suggest the attention-weighted patch means to use Talk2DINO for global tasks such as image-text retrieval.
@@
Considering that in the memory-based model of our framework (that is similar to Li et al. (2023b)) we tackle the modality gap through a projection based on a collection of texts, we tested how much the selection of the texts in the memory bank influences performance. In Table 6, we also report the results obtained by that model when in its memory bank there are also ground-truth captions of the test set (rows marked with GT Memory). This provides a sort of upper bound to performance when varying the text collection used as memory. We observe that in this configuration, the performance only slightly improves (+0.5%), indicating that the model is robust to the choice of the memory bank.

### v3

@@ 5 Conclusions
We introduced a zero-shot framework that shifts from an image-centric to a patch-centric approach, enabling captioning for individual patches and arbitrary aggregations without any region or image supervision. We rely on the strong spatial awareness of DINOv2, whose local image patches have been bridged with the text modality, and we disentangle the training of the decoder network. This flexible and scalable method sets the new state of the art on various regional captioning tasks, including dense and region-based captioning, as well as our newly proposed trace captioning. Results show that our patch-centric approach can effectively bridge the gap between local and global understanding in image captioning, providing a unified framework for multi-granularity captioning tasks in a zero-shot setting. Moreover, our models require a single backbone forward pass to caption multiple regions, facilitating viability in interactive applications.
@@
Despite strong zero-shot performance, our model still lags behind fully supervised, task-specific approaches. The contextual scope of each patch is fixed by the backbone and is not adjusted to meet the user’s intents. Also, the modality gap introduces noise that can cause hallucinations. Future work will focus on improving patch-level semantics, e.g., through image-level captioning loss in weakly-supervised settings, or refine the patch-to-text projection to reduce the modality gap in zero-shot settings.
@@ 6 Implementation Details
For the training of the textual decoder of the memory-based configuration $\phi$, we adopt a prefix GPT2-style decoder-only Transformer with 4 attention heads and 4 layers, following the architecture used by [33]. We train the model on captions from the COCO training set, which also serves as the memory bank $M$ for the projection mechanism, comprising approximately 500k texts. We set the hyperparameters of the projection mechanism as in DeCap ($\tau=0.01$), and use the AdamW optimizer with a weight decay of 0.01. Training proceeds for 10 epochs with a learning rate of $10^{-5}$ and a batch size of 64. A comprehensive overview of the framework operating in the DeCap setting (with the projection mechanism as modality gap mitigation strategy) is provided in Figure 4.
For the external knowledge-based captioning models we performed a training of 15 epochs with a batch size of 80 captions on the same GPT2-style textual decoder using a learning rate of $2\times 10^{-5}$ and a gaussian noise variance of $16\times 10^{-3}$ to replicate the experimental settings of [15] and [68].
@@ 7 Backbone Details
  • CLIP [50]: A foundational model that learns a shared embedding space for images and text through contrastive learning. While being the most used model for global image-text alignment, its patch tokens are known to lack strong spatial and fine-grained semantic information. The input resolution of its training is 224 pixel.
@@
  • DenseCLIP [53]: A fine-tuned version of CLIP that incorporates a pixel-text matching loss to enhance the model’s ability to understand local regions. The official implementation input resolution is 640 pixel for the ViT-B/16 version.
@@
  • INViTE [8]: This method modifies CLIP’s vision transformer to bring patch tokens in the text space by disabling the self-attention mechanism. It employs the same visual encoder of CLIP, trained at 224 pixel input resolution.
@@
  • ProxyCLIP [29]: A model that leverages the local understanding of a DINO backbone to improve CLIP’s patch-level representations. It achieves this by replacing the attention maps in CLIP’s final layer with DINO’s attention maps, effectively transferring DINO’s fine-grained spatial awareness to the CLIP embedding space. The DINO ViT-B/8 version was tested with images at 296 pixel resolution, while the DINOv2 ViT-B/14 at 518 pixel.
  • SigLIP2 [60]: A multilingual vision–language model that improves upon CLIP by enhancing both global alignment and dense localization capabilities through refined training objectives and architectural adjustments. We employed the B/16 variant, which uses a ViT with 16-pixel patches and is trained at an input resolution of 512 pixels. To obtain text-aligned semantic patch representations, we follow the authors’ dense feature extraction strategy and apply the MAP (Mean Attention Pooling) head—normally used to produce the global image representation embedding—individually to each patch token.
@@
  • DINO.txt [25]: This model builds upon a frozen DINOv2 backbone, adding learnable transformer blocks on top. It is then trained with a contrastive objective against a text encoder to align both global and patch-level representations with language. The DINOv2 backbone was trained at the resolution of 518 pixel.
@@
  • Talk2DINO [5]: This model creates a bridge between the CLIP and DINOv2 embedding spaces. It trains a projection to map CLIP text embeddings into the DINOv2 patch space, using DINOv2’s highly meaningful attention maps to identify and align with the most relevant patches during training. The DINOv2 backbone was trained at the resolution of 518 pixel.
8 Comparison with LMMs
Captioning Task: (Dataset) Adaptation strategy Trace (COCO) Dense (VG v1.2) Region-Set (COCO Entities) Image (COCO)
C P C P C P C P CLIP-S
LMMs
Llava1.5 One-Vision (4B) Visual Prompting 21.2 75.9 25.3 76.2 72.2 87.3 91.6 91.1 80.1
Qwen2.5 VL (3B) Visual Prompting 17.4 74.0 19.1 74.2 59.6 84.5 77.9 89.9 81.2
Qwen3 VL (4B) Visual Prompting 9.1 70.7 10.8 74.2 10.8 74.2 19.4 84.2 85.3
Llava1.5 One-Vision (4B) Crop 19.7 75.4 15.1 72.7 75.9 86.7 91.6 91.1 80.1
Qwen2.5 VL (3B) Crop 18.4 73.9 11.9 69.2 68.6 85.9 77.9 89.9 81.2
Qwen3 VL (4B) Crop 7.6 69.3 3.5 68.8 20.2 80.4 19.4 84.2 85.3
Patch-ioner (Our Patch-based Framework)
T2D + Mem. (0.21B) 27.9 78.7 31.9 78.8 109.1 87.5 69.2 87.4 72.8
Table 3: Patch-ioner framework vs. LMMs.
Our framework is designed as a zero-shot regional captioner, relying solely on text-only corpora to train the decoder, and operating without any paired image-text or region-text supervision. While most of the Large Multi-modal Models (LMMs) are trained without region-text supervision, they are trained on massive datasets using image-text pairs, making them inherently supervised solutions.
While this difference in training paradigm places LMMs outside the core assumptions of our zero-shot setup, we include a comparison with representative, high-performing LMMs — Llava1.5 OneVision (4B) [1], Qwen2.5 VL  [3] (3B) and Qwen3 VL (4B) — to provide a strong, external supervised benchmark for our framework’s capabilities.
8.1 LMM Adaptation for Regional Tasks
LMMs typically process a whole image and cannot natively process explicit regional coordinates (such as bounding boxes or traces) as additional inputs, unless specifically trained with region-level box or mask annotations. To enable a fair comparison of our regional tasks, we propose two zero-shot adaptation strategies to inject regional information into the LMM.
  • Visual Prompting: The regional annotations (bounding boxes or traces) are drawn over the image, allowing the model to condition its generation on the spatial input. We then ask the model to describe the annotated image.
  • Cropping: The image is cropped to the bounding box encompassing the regional annotation (be it a trace, a box, or a set of boxes). The cropped image is then fed to the LMM, forcing the model to focus its attention solely on the region of interest.
An example of adaptation for each task with the respective prompt is shown in Figure 3.
8.2 Results Analysis
Table 3 presents the comparison results of our Patch-ioner framework against state-of-the-art LMMs across various captioning granularities.
In the finer-level tasks, specifically Dense Captioning (VG v1.2) and the Trace Captioning (COCO), the pronounced performance gap between our framework and LMMs highlights the fundamental inability of LMMs to effectively reason at the precise region level based solely on a visual annotation or a simple crop without sufficient global context. For these granular tasks, the Visual Prompting strategy is generally preferable for LMMs because it preserves the essential context of the entire image. Furthermore, simple cropping is problematic for concave traces as it fails to precisely delineate the intended region of the image.
In tasks involving broader regions, such as Region-Set Captioning (COCO Entities), the Patch-ioner framework achieves better performance despite LMMs showing closer results, highlighting our higher capability to generate captions specifically guided by the set of input regions. For these broader regions, the Cropping adaptation strategy generally leads to better LMM performance than Visual Prompting, suggesting that Region-Set Captioning demands a lower level of context from the uncropped parts of the image with respect to Dense and Trace Captioning.
For the Image Captioning task, which relies on global understanding, LMMs generally demonstrate superior performance compared to our zero-shot captioner baseline. This is largely expected, as LMMs are bigger and extensively trained on massive image-text pairs.
8.3 Comparison Highlights
The comparison with powerful LMM baselines adapted for regional tasks clearly underscores the unique advantages of the Patch-ioner framework in terms of design, data requirements, parameter efficiency, and inference speed.
Designed for Granularity
The fundamental difference lies in the design objective: our Patch-ioner framework is explicitly engineered for multi-granularity, region-level captioning , while LMMs are built primarily for global image understanding. The need for ad-hoc adaptation strategies for LMMs (Visual Prompting or Cropping) inherently limits their precision, resulting in significantly lower performance on fine-grained regional tasks –as shown in Table 3– compared to our method.
Data and Training Efficiency
Unlike LMMs, which are trained on massive paired image-text datasets and are therefore inherently supervised solutions, our approach adheres to a strict zero-shot paradigm. The text decoder is trained solely on text corpora, eliminating the dependency on costly, large-scale image-text supervision for training the generative module.
Parameter and Computational Efficiency
The Patch-ioner framework is dramatically more lightweight and parameter-efficient than LMMs. We provide a compact model (e.g., $0.21\text{B}$ parameters for the Talk2DINO configuration), contrasting sharply with the multi-billion parameter counts typical of LMMs (e.g., $3\text{B}$ to $4\text{B}$ parameters in the tested models).
Inference Speed and Scalability
One of our framework’s most critical advantages is its inference efficiency for multiple regions within a single image. For an image containing multiple annotations (such as in the Dense Captioning task, which can have over a hundred boxes), we require only a single forward pass of the frozen vision backbone to extract all patch features. These pre-computed patch features can then be reused to caption arbitrary regions. Conversely, an LMM adapted through Cropping or Visual Prompting requires a full inference of the vision-language model for every single annotation, leading to dramatically slower performance when scaling to dense or complex regional tasks
Original Image
Trace Captioning Dense Captioning Region-Set Captioning
Refer to caption Refer to caption Refer to caption
Visual Prompting Adaptation
Refer to caption Refer to caption Refer to caption
Provide a one-sentence caption for the provided region of the image marked by the red dots. Do not mention the red dots or the highlighting. Provide a one-sentence caption for the provided region of the image marked by the red bounding box. Do not mention the red bounding box or the highlighting. Provide a one-sentence caption for the provided regions of the image marked by the red bounding boxes. Do not mention the red bounding boxes or the highlighting.
Cropping Adaptation
Refer to caption Refer to caption Refer to caption
Provide a one-sentence caption for the provided image. Provide a one-sentence caption for the provided image. Provide a one-sentence caption for the provided image.
Figure 3: LMM Adaptation Strategies for Zero-Shot Regional Captioning. This figure visualizes the two input-adaptation strategies used to benchmark Large Multimodal Models (LMMs) on Region-level Captioning tasks: Trace, Dense, and Region-Set Captioning. Top Row: The original image corresponding to each task. Middle Row (Visual Prompting): The task region is spatially indicated by superimposed red annotations (traces or bounding boxes). The LMM input includes this visually prompted image along with a detailed instruction (shown below the images) to describe the marked region while suppressing any mention of the annotations. Bottom Row (Cropping): The input is a minimal crop tightly encompassing the annotated region. The LMM is given a general instruction (shown below the images) to describe the presented image, implicitly forcing focus onto the region of interest.
9 Ablations on Text Decoder Architectures
We tested different architectures and sizes for the text decoder network. In particular, we trained GPT-2 small [51], Gemma 3 270M [58], Qwen 3 0.6B [64], LLaMa 3.2 1B [17] and Qwen 3 1.7B [64]. Details on training hyperparameters are provided in §6. The dataset adopted for these trainings is made of the ground truth captions of COCO train Karpathy split. The total number of different textual captions for that split is 566,747. Table 4 reports the scores obtained when varying the decoder in our framework.
Across all tasks, we observe that increasing the capacity of the textual decoder does not translate into consistent improvements. GPT-2 small (124M) performs surprisingly strongly, outperforming larger models on Trace and Region-Set captioning in terms of CIDEr, while maintaining competitive RefPAC-S scores. Larger decoders such as Gemma 3 (270M), Qwen3 (0.6B), and Llama 3.2 (1B) yield similar or slightly worse results, and the largest tested model, Qwen3 1.7B, exhibits the lowest performance on all CIDEr metrics. These findings suggest that, under our training regime and dataset size (COCO train Karpathy split,  566k captions), model capacity is not the limiting factor. Larger decoders tend to overfit more easily and provide little benefit in a setting where the captioning supervision is narrow in distribution and relatively small compared to their pretraining scale. Moreover, since our overall architecture relies on a strong visual encoder capable of providing semantic local representations, the decoder’s primary role is to map visual representations to fluent text; in this context, compact autoregressive models appear sufficiently expressive. Interestingly, RefPAC-S shows minimal variation across decoder sizes, indicating that semantic alignment with the image features (rather than surface-level text quality) is largely preserved even with larger models. However, the trend of decreasing CIDEr with increasing model size suggests that bigger decoders may introduce unnecessary linguistic diversity that hurts match-based metrics. Overall, these results highlight that changing the decoder affects only marginally the performances.
Captioning Task: (Dataset) Trace (COCO) Dense (VG v1.2) Region-Set (COCO Entities) Image (COCO)
Textual Decoder # Parameters C P C P C P C P CLIP-S
GPT2 124M 27.9 78.7 31.9 78.8 109.1 87.5 69.2 87.4 72.8
Gemma 3 270M 23.5 78.4 25.5 78.5 98.7 87.5 61.0 87.2 73.2
Qwen3 0.6B 23.5 78.4 25.0 78.4 98.7 87.5 59.7 87.2 73.4
LLama 3.2 1B 24.2 78.3 26.0 78.4 101.0 87.4 60.3 87.1 73.5
Qwen3 1.7B 22.9 78.2 24.5 78.2 95.6 87.4 57.9 87.1 73.7
Table 4: Training different decoders. CIDEr (C) and RefPAC-S (P) across four captioning tasks. The model adopted is T2D + Memory ($\approx$ DeCap) trained on COCO train Karpathy split.
10 Ablations on Text-only Training Dataset
In this section, we compare the results obtained when using the classical training strategy adopted by the models such as [33, 68, 43] in the zero-shot captioning task — which consists of training for 10 epochs on the collection of texts from the COCO dataset — with training on a different dataset, to assess generalization capabilities of our framework.
For this comparison, we took the best Patch-ioner configuration — which consists of using Talk2DINO [5] as visual backbone and a memory bank of texts borrowed from COCO as in [33]— and we trained it on a subset of ReLaion.
10.1 Selected dataset: ReLaion 600M
ReLaion is a large-scale image-text dataset containing approximately 600 million image-caption pairs sourced from LAION-2B [55]. Each entry includes multiple machine-generated captions of varying quality, along with a “best caption” field obtained by ranking the alternatives using a CLIP-based scoring function. In our ablation we exclusively use this “best caption” attribute, since it provides a higher-quality textual signal while avoiding noisy or inconsistent descriptions that typically arise in large crawled datasets.
To ensure a fair comparison with the standard COCO-based training strategy from prior work [33, 68, 43], we adopt a subset of ReLaion sized so that one training epoch over that matches the number of training steps of the standard COCO training lasting 10 epochs. Since the COCO training split contains roughly 560k captions, one epoch on a 5.6M-sample subset of ReLaion results in approximately the same number of optimization steps, while a 28.3M-sample subset corresponds to five times more steps. In all experiments, we maintain the same Patch-ioner configuration (our best-performing model) and apply identical optimization settings.
This setup allows us to assess the generalization capability of our framework when trained on broader, noisier web-scale text compared to the highly curated COCO captions that are commonly used in zero-shot captioning pipelines.
10.2 Discussion
The results in Table 5 highlight three main trends.
(1) ReLaion training improves cross-dataset generalization.
Training on ReLaion — even for a single epoch — consistently improves performance across captioning tasks compared to the classical COCO-only training. The 5.6M-sample setting yields clear gains on Trace and Dense Captioning, and the larger 28.3M subset further enhances Region-Set and Image Captioning scores. These improvements indicate that Patch-ioner benefits substantially from the wider linguistic and visual variability captured in ReLaion, despite its noisier nature.
(2) Gains are not uniform across tasks.
While the Trace and Dense tasks benefit the most from ReLaion (e.g., +2.4 CIDEr on Trace and +1.7 on Dense when moving from COCO to ReLaion 5.6M), the COCO-based Region-Set task exhibits a more nuanced behavior. The larger ReLaion subset (28.3M) significantly boosts Region-Set CIDEr, suggesting that recognizing localized entities and relations requires broader caption diversity, which becomes available only at larger scale.
(3) Memory Bank from ReLaion is less effective than the one from COCO.
When employing 500k captions randomly sampled from ReLaion as memory bank, performance drops across many tasks. This suggests that using a collection of texts for the projection mechanism sampled from COCO leads the model to provide captions that are closer to the COCO ground-truth ones, particularly from a syntactic standpoint. In fact, we can notice how the larger performance drop is on the CIDEr metric.
Overall, these findings show that Patch-ioner adapts well to large-scale noisy text, outperforming COCO-trained baselines on most tasks, and that the advantages grow with the size of the ReLaion subset. This demonstrates that our framework can leverage broad web-scale text distributions to improve zero-shot captioning, even when training the textual decoder for a single epoch. However, we pick the model trained on COCO as reference to have a fair comparison with existing models.
Captioning Task: (Dataset) Trace (COCO) Dense (VG v1.2) Region-Set (COCO Entities) Image (COCO)
Text-only Training Dataset C P C P C P C P CLIP-S
COCO Train 27.9 78.7 31.9 78.8 109.1 87.5 69.2 87.4 72.8
ReLaion 5.6M 30.3 79.0 34.1 79.0 109.0 87.7 69.3 87.5 72.5
ReLaion 28.3M 29.7 79.0 33.6 78.9 113.5 87.9 70.6 87.7 73.0
Using a Memory Bank of 500k ReLaion Captions
Relaion 5.6M 26.7 78.4 33.8 79.2 86.3 85.3 54.1 85.0 70.1
ReLaion 28.3M 26.8 78.3 33.4 79.0 86.6 85.3 54.6 85.0 70.5
Table 5: Training on different datasets. CIDEr (C) and RefPAC-S (P) across four captioning tasks. The model adopted is T2D + Memory ($\approx$ DeCap) using the GPT2 textual decoder.
@@ 11 Additional Results: Aggregation Strategies, Input Resolution, Text Collection
Figure 4: Patch-level Captioning. Given an input image, we first extract dense patch-level representations using a vision transformer backbone. For a selected patch, we apply the projection-based mechanism introduced by [33] to mitigate the modality gap and align its representation with the text embedding space. Finally, the transformed embedding is fed into a text decoder trained on a text-only corpus, generating a zero-shot caption for the patch.
@@
For patch-based captioning, we followed Talk2DINO [5] and used an input image resolution of 518x518, obtaining 37 14x14 patches per side when using the Talk2DINO backbone. The original DeCap [33], ViECap [15], MeaCap [68] implementations uses the CLIP B/32 backbone with 224x224 input images with 7 patches per side. We also tested with the CLIP B/16 backbone, resulting in 14 patches per side at 224x224 resolution, and with 592x592 input image size, to obtain the same number of patches as in our framework (37 per side).
@@
In Table 9, we report the results of standard zero-shot image captioning. In addition to the already described weighting schemes, we test one additional configuration for our framework that is central patch, where the decoding is applied to the central patch of the image. We can observe that the most effective strategy for the image captioning task is attention. This is coherent with results from [5], where they suggest the attention-weighted patch means to use Talk2DINO for global tasks such as image-text retrieval.
@@
Considering that in the memory-based model of our framework (that is similar to [33]) we tackle the modality gap through a projection based on a collection of texts, we tested how much the selection of the texts in the memory bank influences performance. In Table 9, we also report the results obtained by that model when in its memory bank there are also ground-truth captions of the test set (rows marked with GT Memory). This provides a sort of upper bound to performance when varying the text collection used as memory. We observe that in this configuration, the performance only slightly improves (+0.5%), indicating that the model is robust to the choice of the memory bank.